OpenStudy (anonymous):

the square root of x+8 is = x-4..

zepdrix (zepdrix):

\[\Large\rm \sqrt{x+8}=x-4\]Hey Grace :) Trying to solve this for x?

zepdrix (zepdrix):

Hmm how bout we square each side?\[\Large\rm \left(\sqrt{x+8}\right)^2=\left(x-4\right)^2\]\[\Large\rm x+8=(x-4)^2\]Do you see how that will help us?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

did you get this? there is another way to do it also